# Teaching Large Language Models to Regress Accurate Image Quality Scores Using Score Distribution

### v2

@@ 1 Introduction
Figure 1: Illustration of our distribution-based soft label. Image A is overexposed, while Image B and C are well-exposed with better and similar quality. “exce.” means “excellent”. To train an MLLM as a quality scorer, continuous scores need to be discretized to discrete level tokens as the training label. Q-Align [70] discretizes human-labeled mean score to a one-hot label, causing information loss. We discretize the score distribution, which is approximated as a Gaussian distribution, to obtain a more accurate soft label. Our soft label maintains better inter-image relationships that the Image B is better than A (${\rm A<B}$), and Image B and C share similar quality (${\rm B\approx C}$), while the one-hot label can not.
@@
Image Quality Assessment (IQA) focuses on perceiving the visual quality of images [84, 26, 66, 41], and is widely adopted in various application scenarios, e.g., image compression and transmission [52, 5], processing smartphone-captured photos [17], and evaluating generation or restoration models [30, 99, 12, 10], etc. With the emergence of Multi-modal Large Language Models (MLLMs) [32, 75, 95, 44], MLLM-based IQA begins to attract more research interests [69, 71, 7, 80, 79]. These methods mainly leverage MLLMs to describe image quality using language, recognizing that language better mirrors human expression, and captures the multifaceted nature of IQA tasks [80].
@@
However, for quality scoring, which needs an accurate numerical score to quantify an image’s quality, existing description-centered MLLMs [69, 71, 80, 79] perform less accurately than traditional IQA methods. This hinders their applications on real-world IQA scenarios. Considering MLLMs’ good generalization ability on various downstream tasks and promising performance on quality description, it is possible for MLLMs to score the image quality. Therefore, in this work, we aim to explore how MLLMs can effectively regress accurate quality scores by leveraging their broad training and inherent adaptability.
@@
We find that the principal challenge in training MLLMs to predict accurate quality scores is the distribution gap between MLLMs’ outputs and quality scores. MLLMs generate outputs as discrete tokens, while the quality score is inherently continuous, usually approximated to a Gaussian distribution. Previous approaches, like Q-Align [70], attempte to address the distribution gap by directly discretizing the mean score (i.e., the expectation of estimated score distribution) into a one-hot label, as depicted in Fig. 1. However, such a one-hot discretization method introduces information loss, causing mismatches between the discretized labels and the original continuous scores. For instance, in Fig. 1, Image A is overexposed with a lower mean score than Image B, but their one-hot labels appear identical. Conversely, Image B and C are well-exposed with similar quality scores, yet their discretized labels are different.
@@
Another key challenge is the dataset variation, where different IQA datasets exhibit distinct distributions, making it hard to simply combine them for co-training. Due to this variation, images from different datasets may have nearly the same quality scores (i.e., linearly re-scaled scores) but drastically different perceptual quality [87, 98]. A previous method, UNIQUE [87], has demonstrated that the fidelity loss [63] under the Thurstone’s model [60] can effectively capture the intra-dataset relationships, thus facilitating the joint training of multiple IQA datasets. However, the fidelity loss requires the model to predict an accurate score distribution, which MLLMs trained with the one-hot label cannot provide. Fortunately, MLLMs trained with our distribution-based soft label can predict such an accurate score distribution, allowing the adaptation of the fidelity loss. We apply this fidelity loss to our MLLM-based quality scorer and demonstrate its effectiveness in this new context.
Equipped with these designs, we successfully develop a distribution-based multi-modal image quality assessment model, termed DeQA-Score. DeQA-Score achieves better score regression results than baseline methods on both in-distribution (e.g., 1.3% in PLCC on KonIQ [21]) and out-of-distribution datasets (e.g., 4.6% in PLCC on LIVE-WILD [18]). The advantage of our DeQA-Score becomes more pronounced with the co-training across multiple IQA datasets. Moreover, our method can further predict the distribution of quality scores, which is closely aligned with the distribution derived from extensive human annotations. We hope that our DeQA-Score will inspire more research to explore the potential of quality score distribution.
@@ 2 Related Works
Full-reference IQA methods compute a similarity score as the quality score between a distorted image and a high-quality reference. Classical works rely on human-designed metrics such as structural similarity [66], phase congruency with gradient magnitude [82], image information [53], etc. The rapid advancement of deep learning has also inspired learning-based IQA methods that measure image quality through data-driven training. Pioneered by LPIPS [84] and PieAPP [48], driven by large-scale IQA datasets [84, 24, 31, 52], deep-learning methods [4, 6, 15, 16, 19, 76, 94, 8] have exhibited quite high accuracy in quality score regression.
@@
Non-reference IQA methods directly regress a quality score without a reference image. Initially, hand-crafted natural image statistics are adopted [37, 40, 41, 42, 43, 50, 58]. Subsequently, deep-learning-based methods [25, 34, 46, 55, 56, 93, 96, 88, 78] replace hand-crafted statistics by learning quality priors from human-annotated IQA datasets [21, 17, 31, 30]. Recent works further improve the regression accuracy by introducing multi-dataset co-training strategy [87], multi-scale features [26], CLIP pre-training [64], multi-dimension attention [73], multitask learning [89], and so on.
@@
Multi-modal Large Language Models (MLLMs) incorporate visual modality into Large Language Models (LLMs) [13, 1, 61, 59] to achieve general visual ability by leveraging the broad training and powerful generalization ability of LLMs. Existing MLLMs [3, 14, 32, 44, 67, 74, 77, 83, 85, 95] have demonstrated a general visual ability and can tackle a variety of multi-modality tasks, like visual question answering [20, 35, 36], document understanding [38, 39, 54], image captioning [2, 9, 81], etc. Given MLLMs’ advanced performance in these high-level perception tasks, it is natural to explore how MLLMs perform in low-level perception tasks that are highly related to IQA.
@@
MLLM-based IQA methods leverages MLLMs’ foundational knowledge to achieve better IQA performance or more detailed assessment results [72, 92]. Q-Bench [68, 91] shows that general MLLMs have some low-level perception ability. [97] evaluates various MLLMs on the widely-adopted Two-Alternative Forced Choice (2AFC) task. Q-Instruct [69] enhances the low-level perception ability of MLLMs by introducing a large-scale dataset. Co-Instruct [71] concentrates on the quality comparison among multiple images. DepictQA [80, 79] performs single-image assessment and paired-image comparison in both full-reference and non-reference settings. Q-Ground [7] focuses on the visual quality grounding task, and performs fine-scale detailed visual quality analysis. Some of these works have proposed MLLM-based quality scoring methods, e.g., binary softmax strategy in [68, 91] and pair-wise voting in [80, 79]. Nonetheless, these works principally focus on the quality description ability using languages.
@@
MLLM-based quality scorer. Initially, Q-Bench [68, 91] proposes a binary softmax strategy, enabling MLLMs to generate quality scores by predicting two discrete quality levels (i.e., good or poor). This strategy is adopted by Q-Instruct [69] and Co-Instruct [71]. Compare2Score [98] achieves quality scoring by teaching MLLMs to compare pair-wise images. Inspired by humans’ annotation process, Q-Align [70] discretizes scores to five discrete levels using one-hot label to train MLLMs, resulting in more accurate score regression. Dog-IQA [33] adopts the one-hot label for training-free IQA by incorporating specific standards and local semantic objects. However, as discussed in Sec. 1, one-hot label causes some information loss and can not fully maintain the inter-image relationship.
@@ 3.1 Revisiting One-hot Label
Training with one-hot label. MLLMs only accept discrete tokens as inputs and outputs. Therefore, as shown in Fig. 2a, continuous scores [^1] Unless otherwise specified, the mean quality scores in all datasets will be normalized to $[1,5]$, and the variances will be normalized accordingly. must be discretized into discrete tokens to train the MLLMs. Q-Align [70] adopts a one-hot discretization method. Specifically, as depicted in Fig. 2b, the score range $[1,5]$ is uniformly divided into five intervals. These intervals represent five rating levels, {“bad”, “poor”, “fair”, “good”, “excellent”}, which is the rating standard defined by [51]. In Fig. 2a, the Mean Opinion Score (MOS) of the image, 4.30, fails into the fifth intervals, resulting in the discretized label “excellent”. Accordingly, the response is formulated as “The quality of this image is excellent”. Using the response as the training labels, following [61, 13], MLLMs can be trained with next token prediction loss, i.e., cross-entropy loss between predicted logits and labels.
Figure 2: (a) MLLMs need discrete level tokens for training. Thus, continuous scores must be converted into discrete level tokens as training labels. (b) Comparison between one-hot label [70] and our soft label. One-hot label uniformly divides the score range into 5 intervals to discretize the mean score, i.e., the expectation of estimated score distribution. One-hot label assumes that different levels are identically independent, which is inaccurate. In contrast, we discretize the estimated Gaussian distribution of the score. Our soft label better preserves the relationships between levels, e.g., “fair” is closer to “good” than “excellent”.
[…]
Table 1: Comparison of discretization precision between Q-Align’s one-hot label [70] and our soft label. The precision of MOS discretization is measured by L1 Error / RMSE and PLCC / SRCC. The distance of original distribution and recovered distribution is measured by JS divergence / Wasserstein distance.
Metrics Discretization KonIQ [21] SPAQ [17] KADID [31]
@@ 3.2 Soft Label
Discretization using score distribution. As demonstrated in [87], the quality score of one image, $x$, is typically modeled as a Gaussian distribution, $x\sim\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma^{2}$ are the Mean Opinion Score (MOS) and variance of multiple human annotations, respectively. The distribution is represented by its probability density function, $f(x)$. As shown in the second row of Fig. 2b, we select five central points, $c_{i}\in\{1,2,3,4,5\},i\in\{0,1,2,3,4\}$, as the centers of the five discrete levels. The width of each level region, $d$, is set as 1. The possibility of the score falling into the $i^{\rm th}$ level is denoted as $p_{i}^{raw}$ and is calculated as:
@@
Ideally, the discretization error of the MOS will be 0. However, the possibility of some levels may slightly fall below zero after the linear transformation. We address this by simply clipping these values to zero. The impact of this clipping on accuracy is negligible, as shown in Tab. 1. After clipping, our discretization error for MOS remains around 10×10$\times$10 × to 35×35$\times$35 × smaller than Q-Align’s one-hot label. Also, our soft label maintains a perfect (i.e., 1.000) Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank-order Correlation Coefficient (SRCC) with the original MOS.
@@
KonIQ [21] SPAQ [17] KADID [31]
$\alpha$ 1.036$\pm$0.041 1.033$\pm$0.085 1.118$\pm$0.091
$\beta$0.005$\pm$0.005 -0.005$\pm$0.014 -0.015$\pm$0.010
@@
Figure 4: Framework of our DeQA-Score trained with soft label. For the “$<$level$>$” token, the KL divergence loss is calculated between predicted probabilities and our soft label. For other tokens, common cross-entropy loss for next token prediction is calculated.
@@ 4.1 Training with Our Soft Label
Model architecture. As illustrated in Fig. 4, we follow the widely adopted MLLM architecture [32, 75] to construct our model. First, a vision encoder is adopted to represent the input image to visual tokens. Then, a vision-to-text connector projects the visual tokens into the textual space. Following [75], we use a vision abstractor as the connector. The abstractor can further reduce the number of visual tokens to 64. Finally, the visual and textual tokens are fused together and fed into an LLM for response prediction.
Next token prediction loss for common response tokens. As described in Sec. 3.1 and illustrated in Fig. 4, the constructed response follows the template, “The quality of this image is $<$level$>$”. For common response tokens other than the level token, we compute the next token prediction loss, i.e., the cross-entropy loss $\mathcal{L}_{ce}$, following [61, 45].
@@
$\mathcal{L}_{kl}=\sum\nolimits_{i}p_{i}\log(p_{i}^{pred}/p_{i}).$ (4)

### v3

@@ 1 Introduction
Figure 1: Illustration of our distribution-based soft label. Image A is overexposed, while Image B and C are well-exposed with better and similar quality. “exce.” means “excellent”. To train an MLLM as a quality scorer, continuous scores need to be discretized to discrete level tokens as the training label. Q-Align [qalign] discretizes human-labeled mean score to a one-hot label, causing information loss. We discretize the score distribution, which is approximated as a Gaussian distribution, to obtain a more accurate soft label. Our soft label maintains better inter-image relationships that the Image B is better than A ($\rm A<B$}), and Image B and C share similar quality ($\rm B\approx C$}), while the one-hot label can not.
@@
Image Quality Assessment (IQA) focuses on perceiving the visual quality of images [lpips, musiq, ssim, niqe], and is widely adopted in various application scenarios, e.g., image compression and transmission [live, phocolens], processing smartphone-captured photos [spaq], and evaluating generation or restoration models [agiqa, agenticir, ultrafusion, prompt_dehaze], etc. With the emergence of Multi-modal Large Language Models (MLLMs) [llava, mplugowl2, minigpt4, gpt4v], MLLM-based IQA begins to attract more research interests [qinstruct, coinstruct, qground, depictqa, depictqav2]. These methods mainly leverage MLLMs to describe image quality using language, recognizing that language better mirrors human expression, and captures the multifaceted nature of IQA tasks [depictqa].
@@
However, for quality scoring, which needs an accurate numerical score to quantify an image’s quality, existing description-centered MLLMs [qinstruct, coinstruct, depictqa, depictqav2] perform less accurately than traditional IQA methods. This hinders their applications on real-world IQA scenarios. Considering MLLMs’ good generalization ability on various downstream tasks and promising performance on quality description, it is possible for MLLMs to score the image quality. Therefore, in this work, we aim to explore how MLLMs can effectively regress accurate quality scores by leveraging their broad training and inherent adaptability.
@@
We find that the principal challenge in training MLLMs to predict accurate quality scores is the distribution gap between MLLMs’ outputs and quality scores. MLLMs generate outputs as discrete tokens, while the quality score is inherently continuous, usually approximated to a Gaussian distribution. Previous approaches, like Q-Align [qalign], attempte to address the distribution gap by directly discretizing the mean score (i.e., the expectation of estimated score distribution) into a one-hot label, as depicted in Fig. 1. However, such a one-hot discretization method introduces information loss, causing mismatches between the discretized labels and the original continuous scores. For instance, in Fig. 1, Image A is overexposed with a lower mean score than Image B, but their one-hot labels appear identical. Conversely, Image B and C are well-exposed with similar quality scores, yet their discretized labels are different.
@@
Another key challenge is the dataset variation, where different IQA datasets exhibit distinct distributions, making it hard to simply combine them for co-training. Due to this variation, images from different datasets may have nearly the same quality scores (i.e., linearly re-scaled scores) but drastically different perceptual quality [unique, compare2score]. A previous method, UNIQUE [unique], has demonstrated that the fidelity loss [fidelity] under the Thurstone’s model [thurstone] can effectively capture the intra-dataset relationships, thus facilitating the joint training of multiple IQA datasets. However, the fidelity loss requires the model to predict an accurate score distribution, which MLLMs trained with the one-hot label cannot provide. Fortunately, MLLMs trained with our distribution-based soft label can predict such an accurate score distribution, allowing the adaptation of the fidelity loss. We apply this fidelity loss to our MLLM-based quality scorer and demonstrate its effectiveness in this new context.
Equipped with these designs, we successfully develop a distribution-based multi-modal image quality assessment model, termed DeQA-Score. DeQA-Score achieves better score regression results than baseline methods on both in-distribution (e.g., 1.3% in PLCC on KonIQ [koniq]) and out-of-distribution datasets (e.g., 4.6% in PLCC on LIVE-WILD [livewild]). The advantage of our DeQA-Score becomes more pronounced with the co-training across multiple IQA datasets. Moreover, our method can further predict the distribution of quality scores, which is closely aligned with the distribution derived from extensive human annotations. We hope that our DeQA-Score will inspire more research to explore the potential of quality score distribution.
@@ 2 Related Works
Full-reference IQA methods compute a similarity score as the quality score between a distorted image and a high-quality reference. Classical works rely on human-designed metrics such as structural similarity [ssim], phase congruency with gradient magnitude [fsim], image information [vif], etc. The rapid advancement of deep learning has also inspired learning-based IQA methods that measure image quality through data-driven training. Pioneered by LPIPS [lpips] and PieAPP [pieapp], driven by large-scale IQA datasets [lpips, pipal, kadid, live], deep-learning methods [WaDIQaM, JSPL, dists, A-DISTS, ghildyal2022stlpips, CVRKD, SRIF, afine] have exhibited quite high accuracy in quality score regression.
@@
Non-reference IQA methods directly regress a quality score without a reference image. Initially, hand-crafted natural image statistics are adopted [ma2017learning, brisque, niqe, moorthy2010two, DIIVINE, saad2012blind, tang2011learning]. Subsequently, deep-learning-based methods [CNNIQA, RankIQA, BPSQM, HyperIQA, graphiqa, CKDN, MetaIQA, zhang2022continual, paq2piq] replace hand-crafted statistics by learning quality priors from human-annotated IQA datasets [koniq, spaq, kadid, agiqa]. Recent works further improve the regression accuracy by introducing multi-dataset co-training strategy [unique], multi-scale features [musiq], CLIP pre-training [clipiqa], multi-dimension attention [maniqa], multitask learning [liqe], and so on.
@@
Multi-modal Large Language Models (MLLMs) incorporate visual modality into Large Language Models (LLMs) [vicuna, gpt4, llama, internlm] to achieve general visual ability by leveraging the broad training and powerful generalization ability of LLMs. Existing MLLMs [flamingo, instructblip, llava, gpt4v, vary, mplugowl, lamm, internvlm, llama_adapter, minigpt4] have demonstrated a general visual ability and can tackle a variety of multi-modality tasks, like visual question answering [vqav2, mmbench, scienceqa], document understanding [chartqa, docvqa, textvqa], image captioning [nocaps, cococap, flickr], etc. Given MLLMs’ advanced performance in these high-level perception tasks, it is natural to explore how MLLMs perform in low-level perception tasks that are highly related to IQA.
@@
MLLM-based IQA methods leverages MLLMs’ foundational knowledge to achieve better IQA performance or more detailed assessment results [iqasurvey_tianhe, iqasurvey_zicheng]. Q-Bench [qbench, qbench_plus] shows that general MLLMs have some low-level perception ability. [2afcprompt] evaluates various MLLMs on the widely-adopted Two-Alternative Forced Choice (2AFC) task. Q-Instruct [qinstruct] enhances the low-level perception ability of MLLMs by introducing a large-scale dataset. Co-Instruct [coinstruct] concentrates on the quality comparison among multiple images. DepictQA [depictqa, depictqav2] performs single-image assessment and paired-image comparison in both full-reference and non-reference settings. Q-Ground [qground] focuses on the visual quality grounding task, and performs fine-scale detailed visual quality analysis. Some of these works have proposed MLLM-based quality scoring methods, e.g., binary softmax strategy in [qbench, qbench_plus] and pair-wise voting in [depictqa, depictqav2]. Nonetheless, these works principally focus on the quality description ability using languages.
@@
MLLM-based quality scorer. Initially, Q-Bench [qbench, qbench_plus] proposes a binary softmax strategy, enabling MLLMs to generate quality scores by predicting two discrete quality levels (i.e., good or poor). This strategy is adopted by Q-Instruct [qinstruct] and Co-Instruct [coinstruct]. Compare2Score [compare2score] achieves quality scoring by teaching MLLMs to compare pair-wise images. Inspired by humans’ annotation process, Q-Align [qalign] discretizes scores to five discrete levels using one-hot label to train MLLMs, resulting in more accurate score regression. Dog-IQA [dogiqa] adopts the one-hot label for training-free IQA by incorporating specific standards and local semantic objects. However, as discussed in Sec. 1, one-hot label causes some information loss and can not fully maintain the inter-image relationship.
@@ 3.1 Revisiting One-hot Label
Training with one-hot label. MLLMs only accept discrete tokens as inputs and outputs. Therefore, as shown in Fig. 2a, continuous scores [^1] Unless otherwise specified, the mean quality scores in all datasets will be normalized to $[1,5]$, and the variances will be normalized accordingly. must be discretized into discrete tokens to train the MLLMs. Q-Align [qalign] adopts a one-hot discretization method. Specifically, as depicted in Fig. 2b, the score range $[1,5]$ is uniformly divided into five intervals. These intervals represent five rating levels, {“bad”, “poor”, “fair”, “good”, “excellent”}, which is the rating standard defined by [five_level]. In Fig. 2a, the Mean Opinion Score (MOS) of the image, 4.30, fails into the fifth intervals, resulting in the discretized label “excellent”. Accordingly, the response is formulated as “The quality of this image is excellent”. Using the response as the training labels, following [llama, vicuna], MLLMs can be trained with next token prediction loss, i.e., cross-entropy loss between predicted logits and labels.
Figure 2: (a) MLLMs need discrete level tokens for training. Thus, continuous scores must be converted into discrete level tokens as training labels. (b) Comparison between one-hot label [qalign] and our soft label. One-hot label uniformly divides the score range into 5 intervals to discretize the mean score, i.e., the expectation of estimated score distribution. One-hot label assumes that different levels are identically independent, which is inaccurate. In contrast, we discretize the estimated Gaussian distribution of the score. Our soft label better preserves the relationships between levels, e.g., “fair” is closer to “good” than “excellent”.
Limitations of one-hot label. First, ideal training with cross-entropy loss using one-hot labels assumes that the five levels are independent (i.e., orthogonal in a latent space). However, intuitively, these levels are not truly independent, e.g., “fair” is closer to “good” than “excellent”. Second, one-hot discretization only considers the mean score, which is the expectation of the score distribution, ignoring the distribution’s uncertainty (i.e., variance). Third, near the border of two levels, images with similar quality can be divided into different levels. For example, in Fig. 1, Image B (MOS 3.38) and Image C (MOS 3.49) share similar mean scores, but are differently classified as “fair” and “good”, due to their proximity to the border of these two levels, i.e., 3.4 in Fig. 2b. Finally, one-hot label causes information loss. In Q-Align, the five levels are respectively assigned scores {1, 2, 3, 4, 5}, with which we calculate the discretization precision in Tab. 1. The discretization error is around $10\times$ to $35\times$ larger than our soft label that we will soon introduce, thus the one-hot label keeps lower correlation with MOS.
Table 1: Comparison of discretization precision between Q-Align’s one-hot label [qalign] and our soft label. The precision of MOS discretization is measured by L1 Error / RMSE and PLCC / SRCC. The distance of original distribution and recovered distribution is measured by JS divergence / Wasserstein distance.
Metrics Discretization KonIQ [koniq] SPAQ [spaq] KADID [kadid]
@@ 3.2 Soft Label
Discretization using score distribution. As demonstrated in [unique], the quality score of one image, $x$, is typically modeled as a Gaussian distribution, $x\sim\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma^{2}$ are the Mean Opinion Score (MOS) and variance of multiple human annotations, respectively. The distribution is represented by its probability density function, $f(x)$. As shown in the second row of Fig. 2b, we select five central points, $c_{i}\in\{1,2,3,4,5\},i\in\{0,1,2,3,4\}$, as the centers of the five discrete levels. The width of each level region, $d$, is set as 1. The possibility of the score falling into the $i^{\rm th}$ level is denoted as $p_{i}^{raw}$ and is calculated as:
@@
Ideally, the discretization error of the MOS will be 0. However, the possibility of some levels may slightly fall below zero after the linear transformation. We address this by simply clipping these values to zero. The impact of this clipping on accuracy is negligible, as shown in Tab. 1. After clipping, our discretization error for MOS remains around $10\times$ to $35\times$ smaller than Q-Align’s one-hot label. Also, our soft label maintains a perfect (i.e., 1.000) Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank-order Correlation Coefficient (SRCC) with the original MOS.
@@
KonIQ [koniq] SPAQ [spaq] KADID [kadid]
[…]
@@
Figure 4: Framework of our DeQA-Score trained with soft label. For the “<<level>>” token, the KL divergence loss is calculated between predicted probabilities and our soft label. For other tokens, common cross-entropy loss for next token prediction is calculated.
@@ 4.1 Training with Our Soft Label
Model architecture. As illustrated in Fig. 4, we follow the widely adopted MLLM architecture [llava, mplugowl2] to construct our model. First, a vision encoder is adopted to represent the input image to visual tokens. Then, a vision-to-text connector projects the visual tokens into the textual space. Following [mplugowl2], we use a vision abstractor as the connector. The abstractor can further reduce the number of visual tokens to 64. Finally, the visual and textual tokens are fused together and fed into an LLM for response prediction.
Next token prediction loss for common response tokens. As described in Sec. 3.1 and illustrated in Fig. 4, the constructed response follows the template, “The quality of this image is <<level>>”. For common response tokens other than the level token, we compute the next token prediction loss, i.e., the cross-entropy loss $\mathcal{L}_{ce}$, following [llama, gpt3.5].
@@
$\mathcal{L}_{kl}=-\sum\nolimits_{i}p_{i}\log(p_{i}^{pred}/p_{i}).$ (4)